OpenStudy (anonymous):

what does 1/5 equal to

OpenStudy (anonymous):

a fraction of 1 over 5

OpenStudy (anonymous):

Decimal form? @cowgirl294

OpenStudy (anonymous):

Find the missing numbers to make these fractions equal. 1 5 = 2 = 15 = 4

OpenStudy (anonymous):

can you take a screenshot?

OpenStudy (anonymous):

1/5 = 0.20

OpenStudy (anonymous):

1.5 or 0.5

OpenStudy (anonymous):

1.5 is equal to 1 1/2 not 1/5

OpenStudy (anonymous):

what is 5/8 = /16

OpenStudy (anonymous):

5/8 = 10/16

OpenStudy (anonymous):

its 5 tenths

OpenStudy (anonymous):

10/16=15/

OpenStudy (anonymous):

10/16 = 15/24

OpenStudy (anonymous):

15/24=/32

OpenStudy (anonymous):

just give someone a medal nd close question

OpenStudy (anonymous):

15/24 = 20/32

OpenStudy (anonymous):

just add 5 to numerator and 8 to denominator from last fraction